Given the following demand and cost function, find the prices and output in each Market and the total profit, Assume that there are only two firms in the market.

P1 = 32 – 2Q1

P1 = 22 – Q2

TC = 10 + 2Q1 + Q2


1
Expert's answer
2021-12-22T13:58:08-0500

"TR1=Q1*P1=(32 \u2013 2Q1)Q1"

"TR2=Q2*P2= (22 \u2013 Q2)Q2"

"MR1=32-4Q1"

"MR2= 22 \u2013 2Q2"

"MC=TC=3"

"32-4Q1=3"

"Q1=7.25"

"22 \u2013 2Q2=3"

"Q2=9.5"

"P1=32 \u2013 2Q1=32-2*7.25=17.5"

"P2= 22 \u2013 Q2=22-9.5=12.5"

"TC = 10 + 2Q1 + Q2=10+2*7.25+9.5=34"

"TR=7.25*17.5+9.5*12.5-34=211.625"